Methane gas, CH4(g), can be compressed into a container to be used as ‘cleaner-burning’ fuel. A company compresses 650.0 L of methane gas at 1.4 atm into a small 175.0 L canister. To do this, the methane is cooled from 23.0°C to lower temperature, and the gas is pressurized to 4.5 psi. Calculate the temperature, in °C, that is required for this compression. {/6}


1
Expert's answer
2021-02-09T03:37:52-0500

V1= 650L

V2= 175L

P1= 1.4atm

P2= 4.5psi

Conversion:

1psi = 0.068atm

4.5psi= 0.068 x 4.5 = 0.306atm

T1= 23°C= 297K

Using the general gas equation

P1V1/T1= P2V2/T2

T2= P2V2T1/P1V1 = 0.306x175x297/1.4 x 650

T2= 17.48K

= 17.48-273°C = -255.52°C